Ex 8.1 Class 6 Maths Question 3.
Write each of the following as decimals:
(a) Seven-tenths
(b) Two tens and nine-tenths
(c) Fourteen point six
(d) One hundred and two ones
(e) Six hundred point eight

Grade:12th pass

1 Answers

Pawan Prajapati
3 years ago
(a) Seven-tenths = 7 x 110 = 0.7 (b) Two tens and nine-tenths = 2 tens + 0 ones + 9 tenths = 20 + 910 = 20.9 (c) Fourteen point six = 14.6 (d) One hundred and two ones = 1 x 100 + 2 x 1 = 100 + 2 = 102 (e) Six hundred point eight = 600.8